# A phase transition for the two-dimensional random field Ising/FK-Ising model

### v1

@@ 4.2 Concentration of $F(h,\omega)$
5 Singularity in strong-disorder regime
This section is devoted to the proof of singularity results in Theorem 1.3. It combines a coarse-graining argument with some near-critical results that will be shown in Section 6. The proof idea is quite similar to Section 3.2, however, careful analysis of the random cluster configuration is involved due to the lack of the nice domain Markov property (DMP) of the RFIM.
5.1 Critical-temperature and high-temperature cases
The critical-temperature and high-temperature cases could be solved together, and the rough intuition is that in both cases the random clusters are not likely to grow too large, or equivalently speaking, there will be plenty of dual paths.
Definition 5.1.
  1. (i)
  2. (ii)
  3. (iii)
For any $\mathcal{O}=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n})$, let
$\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\mathcal{O})=\sum_{\omega:\mathcal{O}=\mathrm{Out}(\omega)}\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\omega)$
be the marginal distribution on the outmost close region. We emphasize that the event $\mathcal{O}=\mathrm{Out}(\omega)$ is measurable under $\omega|_{E(\Lambda_{N})\setminus E(\Omega(\mathcal{O}))}$. We further define $\Phi_{p,\mathcal{O}}^{h}$ to be the product measure of $\phi^{\mathrm{f},h}_{p,\Omega_{i}}~{}(i=1,2,\cdots,n)$. More precisely, let $\omega^{\mathcal{O}}\in\{0,1\}^{E(\Omega(\mathcal{O}))}$ be an edge configuration and $\omega_{i}^{\mathcal{O}}=\omega^{\mathcal{O}}|_{\Omega_{i}},$ then we define
$\Phi_{p,\mathcal{O}}^{h}(\omega^{\mathcal{O}})=\prod_{i=1}^{n}\phi_{p,\Omega_{i}}^{\mathrm{f},h}(\omega_{i}^{\mathcal{O}}).$
Then applying DMP, we obtain that for any configuration $\omega$ and $\mathcal{O}=\mathrm{Out}(\omega)$,
Similar to (3.5) and (3.6), we apply (5.1) to obtain that
[…]
This naturally leads to the following definition:
Definition 5.2.
$\phi_{p,\Lambda_{N}}^{\gamma,0}\Big{(}\big{\{}\mathcal{O}\mid\|\Phi^{0}_{p,\mathcal{O}}-\Phi^{\epsilon h}_{p,\mathcal{O}}\|_{\mathrm{TV}}\geq 1-c\exp(-c^{-1}n)\big{\}}\Big{)}\geq 1-c\exp(-c^{-1}n).$ (5.4)
Lemma 5.3.
Proof.
Recall that $\{B_{i}\}_{1\leq i\leq n}$ is a partition of $\Lambda_{N}$ by $2M$-boxes. Applying the RSW theory (see [18] and also e.g. [17, 30, 25] for other remarkable progress on the RSW theory), there exists a constant $C_{1}>0$ such that $\phi^{\gamma,0}_{p,B_{i}}(\Omega_{i}\neq\emptyset)\geq C_{1}$ where $\Omega_{i}$ is the region satisfying Definition 5.1 and $\gamma$ is an arbitrary boundary condition on $\partial_{int}B_{i}$. Thus, we obtain by DMP and Lemma B.6 that
Let $\mathcal{A}$ denote the event that $1-\|\Phi^{0}_{p,\mathcal{O}}-\Phi^{\epsilon h}_{p,\mathcal{O}}\|_{\mathrm{TV}}\leq C_{3}\exp({-}C_{3}^{-1}n)$ with some $C_{3}>0$ to be determined later. Note that $\Phi^{0}_{p,\mathcal{O}}$ and $\Phi^{\epsilon h}_{p,\mathcal{O}}$ are product measures of $\phi^{\mathrm{f},0}_{p,\Omega_{i}}$ and $\phi^{\mathrm{f},\epsilon h}_{p,\Omega_{i}}$ and Propositions 6.1 and 6.4 show that ${\mathbb{E}}\|\phi^{\mathrm{f},0}_{p,\Omega_{i}}-\phi^{\mathrm{f},\epsilon h}_{p,\Omega_{i}}\|_{\mathrm{TV}}\geq C_{4}>0$. Thus, applying Corollary 3.4 gives that for any $\mathcal{O}$ such that $\eta(\mathcal{O})\geq\frac{C_{1}n}{2}$
Note that we also fix $C_{3}$ here to be the constant given by Corollary 3.4.
$\leq 3c_{1}\exp(-c_{1}^{-1}n).$
The desired result comes from Lemma 5.3. ∎
5.2 The low-temperature case
Now we consider the low-temperature case. Let $M<N$ be a fixed integer with order $O(\epsilon^{-1})$ to be defined later. The main difference from Section 5.1 is that (5.5) does not hold since with high probability there does not exist a dual circuit in $\Lambda_{2M}(u_{i})\setminus\Lambda_{M}(u_{i})$. Thus, we need to consider an open circuit instead of a dual circuit, which provides a wired boundary condition instead of a free boundary condition. Further, we emphasize that the second equality in (5.1) fails since the external field outside $\Omega_{i}$ also influences the measure. Nevertheless, the influence of outside external is manageable if the open circuit is connected to a unique large cluster.
We start with the construction of the desired unique large cluster.
Definition 5.4.
An edge configuration $\omega$ is called well-connected if there exists a unique connected cluster in $\omega\mid_{\Lambda}$ such that its diameter is not less than $N/2.$ And we will call this cluster the main cluster in $\omega$ for convenience. Besides, we write $\mathcal{W}\subset\{0,1\}^{\mathbb{E}(\Lambda_{N})}$ for the set of well-connected configurations.
Remark 5.5.
We highlight that in the definition of a well-connected configuration $\omega$, the uniqueness of the main cluster is confined to $\omega\mid_{\Lambda}$. In other words, despite this restriction, there remains the possibility that certain small clusters within $\omega\mid_{\Lambda}$ may be connected to a large cluster in $\omega$ that differs from the main cluster. This clarification is solely intended to enhance readers’ understanding and does not impact the proof in any way.
Definition 5.6.
  1. (i)
    For all $i=1,2,\cdots,n,$ either $\Omega_{i}=\emptyset$ or $\Lambda_{M/2}(u_{i})\subset\Omega_{i}\subset\Lambda_{M}(u_{i})$ such that $\partial_{int}\Omega_{i}$ is an open circuit in $\omega$ and it is connected to the main cluster of $\omega.$
  2. (ii)
    There does not exists a subset $\Lambda_{M/2}(u_{i})\subset\Omega^{\prime}_{i}\subset\Lambda_{M}(u_{i})$ satisfying (i) such that $\Omega_{i}\subsetneqq\Omega^{\prime}_{i}$.
  3. (iii)
    Let $\Gamma=\cup_{i=1}^{n}(B_{i}\setminus\Omega_{i})$, then $\tilde{\omega}=\omega|_{\Gamma}$.
  4. (iv)
    If $\Omega_{i}\neq\emptyset$, then $h$ is $c$-good with respect to $\Omega_{i}$ where the $c$-good external field will be defined in the Definition 5.7 below.
Figure 2: An illustration of the outmost good open region. The dotted squares are $\Lambda_{M}(u_{i})$ where $\Lambda_{2M}(u_{i})=B_{i}$. The red color denotes that $h$ is $c$-good with respect to the region and the orange color denotes that $h$ is not $c$-good with respect to the region.
We use the notation $\mathcal{O}=\mathrm{Out}(\omega,h)$ to denote that $\mathcal{O}=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n},\tilde{\omega})$ is the outmost good open configuration of $\omega$. Note that for $\mathcal{O}=\mathrm{Out}(\omega,h)$, $\omega$ is well-connected if and only if $\tilde{\omega}$ is well-connected. Thus, we get that $\{\omega:\mathrm{Out}(\omega,h)=(\Omega_{1},\Omega_{2},\cdots,\Omega_{n},\tilde{\omega})\}=\tilde{\omega}\otimes\{0,1\}^{E(\cup_{i=1}^{n}\Omega_{i})}$. Let $\eta(\mathcal{O})$ denote the number of nonempty regions $\Omega_{i}$. We define $\Phi_{p,\mathcal{O}}^{\epsilon h}$ to be FK measure $\phi_{p,\Lambda_{N}}^{{\gamma},\epsilon h}(\cdot\mid\tilde{\omega})$ conditioned on $\omega|_{\Gamma}=\tilde{\omega}$. By calculations in (5.3), it suffices to prove a lower bound for $\|\Phi^{\epsilon h}_{p,\mathcal{O}}-\Phi^{0}_{p,\mathcal{O}}\|_{\mathrm{TV}}$ as in (5.2). Unlike Section 5.1, $\Phi^{\epsilon h}_{p,\mathcal{O}}$ is not a product measure of measures supported on $\Omega_{i}$ because the external field outside $\Omega_{i}$ will influence the measure on $\Omega_{i}$. Thus, we give the following definition of good external field.
Definition 5.7.
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
Lemma 5.8.
There exists a constant $\theta_{0}>0$ such that for any $0<\theta<\theta_{0}$ and $\theta M^{-1}\leq\epsilon\leq 2\theta M^{-1}$, we have the following. For any region $\Lambda_{M}\subset\Omega\subset\Lambda_{2M}$ and any $c$-good external field $h$ with respect to $\Omega$, let $X^{\pm}=\sum_{\mathcal{C}\in\mathfrak{C}\setminus\{\mathcal{C}_{*}\}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)\pm\frac{\epsilon h_{\mathcal{C}_{*}}}{T}$. Then we have
$\langle X^{\pm}\rangle^{0}_{p,\Omega}\leq\log\big{(}\langle\exp(X^{\pm})\rangle^{0}_{p,\Omega}\big{)}-c_{1}$
for some $c_{1}>0$ depending only on $\theta$ and $c$.
Definition 5.9.
Lemma 5.10.
Lemma 5.11.
For any $\mathcal{O}$ such that $\eta(\mathcal{O})\geq c_{1}n$, we have
where $c_{2}>0$ does not depend on $\mathcal{O}$ or $\epsilon h$.
Proof.
Fix any external field $\epsilon h$ and $\mathcal{O},$ let $I$ denote the set of indices such that $\Omega_{i}\neq\emptyset$ (thus $h$ is good with respect to $\Omega_{i}$). For any $\mathrm{Out}(\omega,h)=\mathcal{O}$ and $i\in I$, let $\mathfrak{C}^{i}$ be the collection of clusters in $\omega\mid_{\Omega_{i}}$ and let $\mathcal{C}_{*}^{i}$ be the boundary cluster in $\Omega_{i}$. Let
Then we can compute the Radon-Nikodym derivative between $\Phi^{0}_{p,\mathcal{O}}$ and $\Phi^{\epsilon h}_{p,\mathcal{O}}$
$\langle X_{i}^{\pm}\rangle^{0}_{p,\Omega_{i}}\leq\log\left(\langle\exp(X_{i}^{\pm})\rangle^{0}_{p,\Omega_{i}}\right)-C_{1}$
Combining (5.11) with (5.9) shows that for any $\omega\in\mathscr{S}^{+}\cap\mathscr{S}^{-}$
The desired result comes from combining (5.12) with (5.10). ∎
Finally, we come back to the proof of Lemmas 5.8 and 5.10.
Proof of Lemma 5.8.
$-\sum_{\mathcal{C}\in\mathfrak{C}(\omega_{2})\setminus\{\mathcal{C}_{*}(\omega_{2})\}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)-\frac{\epsilon^{2}|\Omega|-|\mathcal{C}_{*}(\omega_{2})|}{2T^{2}}-\Bigg{|}\frac{\epsilon^{2}(\big{|}|\mathcal{C}_{*}(\omega_{1})|-|\mathcal{C}_{*}(\omega_{2})|\big{|})}{2T^{2}}\Bigg{|}$
$\geq\frac{C_{2}}{2}.$
$\langle X^{+}\rangle^{0}_{p,\Omega}\leq\log\big{(}\langle\exp(X^{+})\rangle^{0}_{p,\Omega}\big{)}-C_{5}.$
By symmetry, the same result holds for $X^{-}$. ∎
Proof of Lemma 5.10.
For a box $B_{i}=\Lambda_{2M}(u_{i}),$ let $\mathcal{G}(u)$ denote the event that there exist two open circuits in the annuli $\Lambda_{4M}(u)\setminus\Lambda_{2M}(u)$ and $\Lambda_{M}(u)\setminus\Lambda_{M/2}(u)$ respectively, and these two circuits are connected in $\omega.$ Then by the RSW theory (see [18] for example), there exists a constant $C_{1}>0$ only relies on $p$ such that for any boundary condition $\gamma$ on $\partial_{int}\Lambda_{4M}(u_{i}),$
$\phi^{\gamma,0}_{p,\Lambda_{4M}(u_{i})}(\mathcal{G}(u))\geq 1-\exp(-C_{1}M).$
Therefore, we define $\mathcal{Q}=\left(Q_{x}=\mathbf{1}_{\mathcal{G}(2Mx)}:x\in\Lambda_{\frac{N}{2M}}\right)$ to be a random vertex configuration on the box $\Lambda_{\frac{N}{2M}},$ and then by [28], we get that $\mathcal{Q}$ dominates a Bernoulli site percolation $\mathbb{P}_{p},$ where $p$ can be arbitrary close to 1 as $M$ goes to infinity. Of course under the measure $\mathbb{P}_{p},$ there exists a unique large cluster with diameter at least $\frac{N}{4M}$ and volume at least $\frac{C_{2}N^{2}}{M^{2}}$ with probability $1-\exp(-C_{3}\frac{N}{M}).$ Thus, we get
$\phi^{\gamma,0}_{p,\Lambda_{N}}(\omega\text{ is well-connected})\geq 1-\exp\left(-C_{4}\frac{N}{M}\right)$
by noticing that a connected component in $\mathcal{Q}$ induces an open cluster in $\omega.$
We choose $\theta=1$ and $M$ such that $\theta M^{-1}\leq\epsilon\leq 2\theta M^{-1}$. Futhermore, we choose $c$ in Definition 5.7 be the first constant in Lemma 6.14. Then by the results in Section 6.3 (see Lemmas 6.12, 6.13 and 6.14), we get that the external field on the region $\Omega$ is good with a positive probability. Hence by Lemma B.6, we get that
and the desired result comes from Markov’s inequality. ∎
@@ 6 Near-critical behaviour
As shown in Section 5, the core of proving singularity is to lower-bound the influence of external field on an $M$ box (with $M\sim\epsilon^{-\frac{1}{\alpha(T)}}$) and we will do it in this section for all temperature regimes. To be precise, we will show that as long as $\epsilon>cN^{-\alpha(T)}$ for some constant $c>0$, ${\mathbb{E}}\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{\gamma,\epsilon h}\|_{\mathrm{TV}}$ is bounded away from 00 for $T\geq T_{c}$ in Sections 6.1 and 6.2. A similar lower bound in the low-temperature regime will be obtained in Section 6.3 with different emphasis on the ingredients used in Section 5.2. Finally, as a byproduct, we will show that $\alpha(T)$ is the critical threshold of the choice of exponent by proving the upper bound for total variation distance on an $M$ box in all temperature regimes in Section 6.4.
@@ 6.1 The critical-temperature case
There exist constants $c_{1},c_{2}$ depending only on $\theta$ such that for any disorder strength $\theta N^{-\frac{15}{16}}\leq\epsilon\leq 2\theta N^{-\frac{15}{16}}$, domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$ and boundary condition $\gamma$ we have $\mathbb{P}(\mathcal{H}_{\mathrm{cri}}(c_{1}))\geq c_{2}$.
@@
where $X$ denotes a normal random variable with mean 00 and variance $1$. Combining (6.2) and (6.3) shows that
@@
Combining Lemmas 4.3 and 6.3 and integrating over $\mathcal{H}^{0}\cap\mathcal{H}_{\mathrm{cri}}(c_{1})$, we obtain that
@@ 6.2 The high-temperature case
In this subsection, we consider the high-temperature case and fix $p<p_{c}$.
@@
Fix $T>T_{c}(2)$ and thus $p<p_{c}$. For any constant $\theta>0$, there exists a constant $c=c(\theta,p)>0$ such that ${\mathbb{E}}\|\phi_{p,\Omega}^{\mathrm{f},\epsilon h}-\phi_{p,\Omega}^{\mathrm{f},0}\|_{\mathrm{TV}}\geq c$ for any disorder strength $\theta N^{-\frac{1}{2}}\leq\epsilon\leq 2\theta N^{-\frac{1}{2}}$.
@@
In fact, the proof presented in this subsection is valid only for values of $\theta\in(0,\theta_{0})$ with $\theta_{0}$ small enough. Regarding the comprehensive proof, Lemma 5.3 establishes a coarse-graining framework that generalizes to all positive $\theta$, using the case that $\theta\in(0,\theta_{0})$ as an input. While this arrangement may seem somewhat confusing, particularly for readers focusing solely on the proof of Proposition 6.4, it is more appropriately integrated within the overall structure and organization of the paper.
@@
Let $\mathcal{F}\subset\{0,1\}^{\mathrm{E}(\Omega)}$ denote the set of edge configurations that $\max_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|\leq N^{0.1}$ and $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{4}\leq c_{1}N^{2}$ where $c_{1}$ comes from Lemma C.4. We use the notation $\mathcal{H}^{0}_{\mathrm{hig}}$ to denote the set of external field such that
@@
Here and in the following, we write
$\left\langle X\mid\mathcal{F}\right\rangle:=\frac{\left\langle X\cdot\mathbf{1}_{\mathcal{F}}\right\rangle_{p,\Omega}^{\mathrm{f},0}}{\left\langle\mathbf{1}_{\mathcal{F}}\right\rangle_{p,\Omega}^{\mathrm{f},0}}$
for any variable $X$ and event $\mathcal{F}$.
@@
There exists constant $c>0$ depending only on $p$ such that for any disorder strength $\epsilon>0$ and domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$, we have
@@
With Lemmas 6.8, 6.9 and 6.10, we are ready to prove Proposition 6.4.
@@
To show the lower bound, we consider $h\in\mathcal{H}^{0}_{\mathrm{hig}}\cap\mathcal{H}_{\mathrm{hig}}^{1}(c_{1})\cap\mathcal{H}_{\mathrm{hig}}^{2}(c_{3}\theta^{4})$ where $c_{1}$ and $c_{3}$ are the constants defined in Lemmas 6.9 and 6.10. Applying Lemmas 6.8, 6.9 and 6.10, it suffices to show that $\|\phi_{p,\Omega}^{\mathrm{f},\epsilon h}-\phi_{p,\Omega}^{\mathrm{f},0}\|_{\mathrm{TV}}\geq C$ for any $h\in\mathcal{H}_{\mathrm{hig}}^{0}\cap\mathcal{H}_{\mathrm{hig}}^{1}(c_{1})\cap\mathcal{H}_{\mathrm{hig}}^{2}(c_{3}\theta^{4})$ and some constant $C>0$ independent of $h$. Let $X(\omega)=\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}-\frac{\epsilon^{2}|\Omega|}{2T^{2}}$. Applying (6.4), we have
@@
$\Bigg{|}\log\left(\frac{\phi_{p,\Omega}^{\mathrm{f},\epsilon h}(\omega)}{\phi_{p,\Omega}^{\mathrm{f},0}(\omega)}\right)\Bigg{|}=\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)+\log\left(\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p,\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p,\Omega}}\right)\Bigg{|}$
$\geq-\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)-\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}+\Bigg{|}X(\omega)+\log\left(\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p,\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p,\Omega}}\right)+\frac{\epsilon^{2}|\Omega|}{2T^{2}}\Bigg{|}.$
Let $A=-\log\left(\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p,\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p,\Omega}}\right)-\frac{\epsilon^{2}|\Omega|}{2T^{2}}$, then for any $\omega\in\mathscr{S}_{1}(A)\cap\mathscr{S}_{2}$, we have
$\Bigg{|}\log\left(\frac{\phi_{p,\Omega}^{\mathrm{f},\epsilon h}(\omega)}{\phi^{\mathrm{f},0}_{p,\Omega}(\omega)}\right)\Bigg{|}\geq\frac{\sqrt{c_{3}}\theta^{2}}{4}-c_{1}\epsilon^{4}N^{2}\geq\frac{\sqrt{c_{3}}\theta^{2}}{4}-16c_{1}\theta^{4}.$
@@
Let $\mathcal{B}\subset\mathbb{R}^{\Omega}\otimes\{0,1\}^{\mathrm{E}(\Omega)}$ denote the event that $\sum_{\mathcal{C}\in\mathfrak{C}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\geq-C_{1}\epsilon^{4}N^{2}$ where $C_{1}>0$ is some constant to be determined. Since we have $f(x)\geq\frac{x^{2}-x^{4}}{2}$, thus we obtain for any $\omega\in\mathcal{F}$,
@@
Integrating over $\omega\in\mathcal{F}$ and applying Lemma C.4 shows that
$\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{f},0}(\mathcal{B}^{c})\leq C_{5}\exp\left(-C_{5}^{-1}N\right)+\frac{C_{4}}{N^{2}}.$
The desired result on $\mathbb{P}(\mathcal{H}_{\mathrm{hig}}^{1}(C_{1}))$ thus comes from the Markov inequality and letting $N$ big enough. ∎
@@
Note that for $\omega\in\mathcal{F}$, we have the bound
$\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}-\sum_{\mathcal{C}\in\mathfrak{C},\tilde{\mathcal{C}}\in\tilde{\mathfrak{C}}}|\mathcal{C}\cap\tilde{\mathcal{C}}|^{2}\leq\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\leq(2N)^{4}.$
Combined with (6.12), it yields that
@@
Recall the definition of $\mathcal{F}$ in Definition 6.6 , we get for any $\omega\in\mathcal{F}$ that
$\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\leq\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{4}\leq cN^{2}.$
Combined with (6.15), it yields that

### v2

@@ 5 Near-critical behaviour
The crucial input for Section 6 is to lower-bound the influence of external field on an $M$ box (with $M\sim\epsilon^{-\frac{1}{\alpha(T)}}$) and we will do it in this section for all temperature regimes. To be precise, we will show that as long as $\epsilon>cN^{-\alpha(T)}$ for some constant $c>0$, ${\mathbb{E}}\|\phi_{p,\Lambda_{N}}^{\gamma,0}-\phi_{p,\Lambda_{N}}^{\gamma,\epsilon h}\|_{\mathrm{TV}}$ is bounded away from 00 for $T\geq T_{c}$ in Sections 5.1 and 5.2. A similar lower bound in the low-temperature regime will be obtained in Section 5.3 with different emphasis according to the ingredients that will be used in Section 6.2. Finally, as a byproduct, we will show that $\alpha(T)$ is the critical threshold of the choice of exponent by proving the upper bound for total variation distance on an $M$ box in all temperature regimes in Section 5.4.
We point out that the primary computational effort of this work is concentrated in this section. Readers less interested in technical derivations may proceed directly to Section 6, as the continuity of the exposition remains intact when assuming the core results from Section 5.
@@ 5.1 The critical-temperature case
There exist constants $c_{1},c_{2}>0$ depending only on $\theta$ such that for any disorder strength $\theta N^{-\frac{15}{16}}\leq\epsilon\leq 2\theta N^{-\frac{15}{16}}$, domain $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$ and boundary condition $\gamma$ we have $\mathbb{P}(\mathcal{H}_{\mathrm{cri}}(c_{1}))\geq c_{2}$.
@@
where $X$ denotes a normal random variable with mean 00 and variance $1$ with law $\mathbb{P}_{X}$. Combining (5.2) and (5.3) shows that
@@
Combining Lemmas 4.3 and 5.3 and integrating over $\mathcal{H}^{0}\cap\mathcal{H}_{\mathrm{cri}}(c_{1})$, we obtain that
@@ 5.2 The high-temperature case
In this subsection, we consider the high-temperature case.
@@
Fix $T>T_{c}(2)$ and thus $p<p_{c}$. Then there exists a constant $\theta_{0}=\theta_{0}(p)>0$ such that for any constant $\theta\in(0,\theta_{0})$, there exists a constant $c=c(\theta,p)>0$ such that ${\mathbb{E}}\|\phi_{p,\Omega}^{\mathrm{f},\epsilon h}-\phi_{p,\Omega}^{\mathrm{f},0}\|_{\mathrm{TV}}\geq c$ for any disorder strength $\theta N^{-\frac{1}{2}}\leq\epsilon\leq 2\theta N^{-\frac{1}{2}}$.
@@
One could easily remove the restriction of $\theta\in(0,\theta_{0})$ via a coarse graining method, see Section 6.1 for a detailed coarse graining framework (but with a different purpose).
@@
Let $\mathcal{F}\subset\{0,1\}^{\mathrm{E}(\Omega)}$ denote the set of edge configurations that $\max_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|\leq N^{0.1}$ and $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{4}\leq 2c_{1}N^{2}$ where $c_{1}$ comes from Lemma C.4. We use the notation $\mathcal{H}^{0}_{\mathrm{hig}}$ to denote the set of external field such that
@@
Here and in the following, we write $\left\langle X\mid\mathcal{F}\right\rangle:=\frac{\left\langle X\cdot\mathbf{1}_{\mathcal{F}}\right\rangle_{p,\Omega}^{\mathrm{f},0}}{\left\langle\mathbf{1}_{\mathcal{F}}\right\rangle_{p,\Omega}^{\mathrm{f},0}}$ for any variable $X$ and event $\mathcal{F}$.
@@
With Lemmas 5.8, 5.9 and 5.10, we are ready to prove Proposition 5.4.
@@
To show the lower bound, we consider $h\in\mathcal{H}^{0}_{\mathrm{hig}}\cap\mathcal{H}_{\mathrm{hig}}^{1}(c_{1})\cap\mathcal{H}_{\mathrm{hig}}^{2}(c_{3}\theta^{4})$ where $c_{1}$ and $c_{3}$ are the constants defined in Lemmas 5.9 and 5.10. Applying Lemmas 5.8, 5.9 and 5.10, it suffices to show that $\|\phi_{p,\Omega}^{\mathrm{f},\epsilon h}-\phi_{p,\Omega}^{\mathrm{f},0}\|_{\mathrm{TV}}\geq C$ for any $h\in\mathcal{H}_{\mathrm{hig}}^{0}\cap\mathcal{H}_{\mathrm{hig}}^{1}(c_{1})\cap\mathcal{H}_{\mathrm{hig}}^{2}(c_{3}\theta^{4})$ and some constant $C>0$ independent of $h$. Let $X(\omega)=\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}-\frac{\epsilon^{2}|\Omega|}{2T^{2}}$. Applying (5.4), we have
@@
$\Bigg{|}\ln\left(\frac{\phi_{p,\Omega}^{\mathrm{f},\epsilon h}(\omega)}{\phi_{p,\Omega}^{\mathrm{f},0}(\omega)}\right)\Bigg{|}=\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)+\ln\left(\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p,\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p,\Omega}}\right)\Bigg{|}$
$\geq-\Bigg{|}\sum_{\mathcal{C}\in\mathfrak{C}}f\left(\frac{\epsilon h_{\mathcal{C}}}{T}\right)-\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\Bigg{|}+\Bigg{|}X(\omega)+\ln\left(\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p,\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p,\Omega}}\right)+\frac{\epsilon^{2}|\Omega|}{2T^{2}}\Bigg{|}.$
Let $A=-\ln\left(\frac{\mathcal{Z}^{\mathrm{f},0}_{\phi,p,\Omega}}{\mathcal{Z}^{\mathrm{f},\epsilon h}_{\phi,p,\Omega}}\right)-\frac{\epsilon^{2}|\Omega|}{2T^{2}}$, then for any $\omega\in\mathscr{S}_{1}(A)\cap\mathscr{S}_{2}$, we have
$\Bigg{|}\ln\left(\frac{\phi_{p,\Omega}^{\mathrm{f},\epsilon h}(\omega)}{\phi^{\mathrm{f},0}_{p,\Omega}(\omega)}\right)\Bigg{|}\geq\frac{\sqrt{c_{3}}\theta^{2}}{4}-c_{1}\epsilon^{4}N^{2}\geq\frac{\sqrt{c_{3}}\theta^{2}}{4}-16c_{1}\theta^{4}.$
@@
Let $\mathcal{B}\subset\mathbb{R}^{\Omega}\otimes\{0,1\}^{\mathrm{E}(\Omega)}$ denote the event that $\sum_{\mathcal{C}\in\mathfrak{C}}f(\frac{\epsilon h_{\mathcal{C}}}{T})-\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}\geq-C_{1}\epsilon^{4}N^{2}$ where $C_{1}>0$ is some constant to be determined. Since $f(x)\geq\frac{x^{2}-x^{4}}{2}$, we obtain that for any $\omega\in\mathcal{F}$,
@@
Integrating over $\omega\in\mathcal{F}$ and applying Lemma C.4 shows that $\mathbb{P}\otimes\phi_{p,\Omega}^{\mathrm{f},0}(\mathcal{B}^{c})\leq C_{5}\exp\left(-C_{5}^{-1}N^{0.05}\right)+\frac{C_{4}}{N^{2}}.$ The desired result on $\mathbb{P}(\mathcal{H}_{\mathrm{hig}}^{1}(C_{1}))$ thus comes from the Markov inequality and letting $N$ big enough. ∎
@@
Note that for $\omega\in\mathcal{F}^{c}$, we have the bound $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}-\sum_{\mathcal{C}\in\mathfrak{C},\tilde{\mathcal{C}}\in\tilde{\mathfrak{C}}}|\mathcal{C}\cap\tilde{\mathcal{C}}|^{2}\leq\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\leq(2N)^{4}.$ Combined with (5.12), it yields that
@@
Recall the definition of $\mathcal{F}$ in Definition 5.6 , we get for any $\omega\in\mathcal{F}$ that $\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{2}\leq\sum_{\mathcal{C}\in\mathfrak{C}}|\mathcal{C}|^{4}\\ \leq cN^{2}.$ Combined with (5.15), it yields that